OpenStudy (anonymous):

-9y-12(2-3y)=-25-2(y+6)

OpenStudy (anonymous):

What is the context its in?

OpenStudy (zzr0ck3r):

distribute on both sides -9y -24+36y = -25-2y-12 now solve for y, can you do this?

OpenStudy (zzr0ck3r):

get all the y's on one side -9y+36y+2y = 24-25-12 can you do this? @countrygirlmadeover

OpenStudy (anonymous):

i was looking at it but no i dont know how

OpenStudy (zzr0ck3r):

what is -9y+36y+2y? what is 24-25-12?

OpenStudy (anonymous):

i subtract all them lol

OpenStudy (zzr0ck3r):

do you not know how to do -9y+36+2y?

OpenStudy (anonymous):

-9y+36y+2y = 24-25-12 this is 29y =-13 from here u can take on by dividing both sides by 29

OpenStudy (anonymous):

no i dont know anything about algbra

OpenStudy (zzr0ck3r):

you might want to back up a little then, google combing like variables.

OpenStudy (anonymous):

i try but i just dont understand it i need a teacher

OpenStudy (anonymous):

what do i divide by29

OpenStudy (anonymous):

what do i do to get my answer?